# The Relative Entropy of Expectation and Price

### v3

@@ I Introduction
The objective is to create foundations for pricing that do not rely on unrealistic assumptions on market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing this objective, the entropic risk metric has dual application: as the determinant of margin used to protect against default; and as the a priori measure of performance used to optimise the investment strategy.
@@ III Entropic pricing and hedging
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The case studies in Figures 7 and 8 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer spread around the mid model that prices the residual risk after hedging. To further develop intuition on the meaning of the hedge and price conditions and their solutions, it is useful to consider boundary cases where they simplify.
@@ III.4 Markovian diffusion
With this assumption, the price model is established on the assumed divisible process for the state vector $x$ and the underlying price vector $q[t,x]$ given as a function of time and state. Using the expression for Itô’s lemma in the price measure, the calibration, hedge and price conditions are transformed into partial differential-integral form. The calibration condition becomes:
@@ III.5 Heston model
The funding price has deterministic increment with mean $br\,dt$ where $r$ is the funding rate. The underlying price has increment with mean $q\mu\,dt$ and variance $q^{2}\nu\,dt$ where $\mu$ is the mean rate. The variance has increment with mean $(\bar{\nu}-\nu)\kappa\,dt$ and variance $\nu\sigma^{2}\,dt$ where $\kappa$ is the mean reversion rate, $\bar{\nu}$ is the mean reversion level and $\sigma$ is the volatility of variance. The final parameter $\rho$ is the correlation between the increments of the underlying price and the variance.
Since the funding security has deterministic return, the implied funding rate is $r$ and the implied funding spread is $s=0$. In order to satisfy the self-funding condition, the portfolio with underlying weight $\phi$ must have funding weight $-\phi q/b$, and the cumulant generating function in the price measure $\mathbb{E}^{\phi}$ satisfies:
@@ III.6 Data-driven model
For this data-driven approach, the expectation measure is represented as a uniform distribution on a finite set of states, which could for example be sampled from a generative model trained on historical market data. The calibration condition is solved for the unit optimal portfolio $\phi$ via the optimisation problem:
@@
The price measure is then supported on the same set of states but with non-uniform probabilities:
The hedge and price conditions are solved simultaneously for the hedge portfolio $\delta$ and the initial derivative price $p$ via the optimisation problem:
@@
These expressions are amenable to machine learning methods and are the starting point for a variety of applications in deep hedging.
This example demonstrates the reach of the entropic framework. With no theory to support development beyond the information that the data expresses, data-driven models need a more general foundation that does not depend on unrealistic assumptions such as complete markets, continuous hedging or risk-free funding. Advocates of deep hedging address this by replacing dynamic risk neutralisation with dynamic risk optimisation, using convex risk metrics to assess strategic performance. While they suffer from lack of explainability, data-driven models can uncover complex relationships between economic variables that may be hard to capture theoretically.
@@ V Entropic XVA
Figure 13: In this example, the trader receives the initial derivative price $p$ and deploys additional capital $c$ to guarantee settlement, with $c+p\geq 0$. These proceeds are deposited in a margin account with unit price $b>0$. The final derivative price $p+dp$ is settled up to the value $(c+p)(1+db/b)$ of the margin account, with any remainder returned to the trader. Default occurs when the margin account is insufficient to meet the contractual obligation. In the expression for the final settlement, the cap is denoted by the superscript.
@@
Figure 14: Two counterparties deploy capital $c_{1}$ and $c_{2}$ respectively as their guarantee of settlement, with $c_{1}-p\geq 0$ and $c_{2}+p\geq 0$, and they default on their contractual obligation if this is insufficient. Net of all considerations, the counterparties share the combined proceeds from the margin accounts, and the final return on capital is floored and capped by the respective default events. In the expressions for the final settlement and capital, the floor is denoted by the subscript and the cap is denoted by the superscript.
@@
Figure 15: Eliminating internal settlements, the trade is simplified as a funding pool with contributions from both counterparties shared according to the contractual settlement of the derivative security adjusted for default. Using the entropic margin model, these settlements can be encoded in a smart contract for decentralised distribution.
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a trade with initial price $p$. Excess capital is deposited in the margin account with unit price $b>0$. The trade is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
As shown in Figures 15, 15 and 15, net of all considerations the capital returned from the trade is:
$c_{1}+dc_{1}$ $=(b+db)\!\left(\frac{c_{1}}{b}+d\frac{p}{b}\right)_{0}^{(c_{1}+c_{2})/b}$ (84)
$=c_{1}(1+db/b)+\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
$c_{2}+dc_{2}$ $=(b+db)\!\left(\frac{c_{2}}{b}-d\frac{p}{b}\right)_{0}^{(c_{1}+c_{2})/b}$
$=c_{2}(1+db/b)-\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
where the subscript denotes the floor and the superscript denotes the cap. The combined proceeds from the margin accounts are shared between the counterparties, adjusting for the funded return on the derivative security, floored and capped so that the capital is positive.
The first counterparty defaults when the floor is breached and the second counterparty defaults when the cap is breached. Each counterparty assesses the threshold for default using their expectation measure $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$, and default is avoided when the capital exceeds the maximum liability:
$c_{1}$ $\geq p-\operatorname*{ess\,inf}\nolimits_{1}\!\left[\frac{P}{1+db/b}\right]$ (85)
$c_{2}$ $\geq\operatorname*{ess\,sup}\nolimits_{2}\!\left[\frac{P}{1+db/b}\right]-p$
where $P:=p+dp$ is the final price. Setting these inequalities to equalities generates the minimum capital required to fully guarantee settlement, under the assumptions of the respective expectation measures. More generally, the counterparties can exchange certainty of settlement for a reduction in capital. In the entropic margin model, this is parametrised by the default aversion $\beta_{1}$ and $\beta_{2}$ with:
$m_{1}$ $:=c_{1}-p=-\mathbb{E}_{1}[\beta_{1}]\!\left[\frac{P}{1+db/b}\right]$ (86)
$m_{2}$ $:=c_{2}+p=\mathbb{E}_{2}[-\beta_{2}]\!\left[\frac{P}{1+db/b}\right]$
where the case $\beta_{1}=\infty$ means the first counterparty provides full default protection and the case $\beta_{2}=\infty$ means the second counterparty provides full default protection.
Viability of the trade is assessed by the counterparty using their risk aversion $\alpha_{1}$ and $\alpha_{2}$. Consider the threshold for execution given by the condition:
$\mathbb{E}_{1}[\alpha_{1}]\!\left[\frac{c_{1}+dc_{1}}{1+db/b}\right]$ $\geq c_{1}$ (87)
$\mathbb{E}_{2}[\alpha_{2}]\!\left[\frac{c_{2}+dc_{2}}{1+db/b}\right]$ $\geq c_{2}$
on the entropy-adjusted mean return on funded capital. These conditions impose bounds on the viable price:
$p$ $\leq p_{1}:=\mathbb{E}_{1}[\alpha_{1}]\!\left[\left(\frac{P}{1+db/b}\right)_{\!-m_{1}}^{\!m_{2}}\right]$ (88)
$p$ $\geq p_{2}:=\mathbb{E}_{2}[-\alpha_{2}]\!\left[\left(\frac{P}{1+db/b}\right)_{\!-m_{1}}^{\!m_{2}}\right]$
The trade is viable when the initial price $p$ is in the range $p_{2}\leq p\leq p_{1}$, and is therefore not viable when $p_{1}<p_{2}$. Price is thus constrained to the sector bounded by these viability thresholds, equilibrating to a price within the range. Since the entropy-adjusted mean is monotonic as a function of the notional, the range narrows as the notional increases and there is an equilibrium point at which no further trades are viable. Equilibrium is thus discovered as the solution of the relationship:
[…]
for the initial price per unit notional $p$ and the equilibrium notional $\lambda$. Terminating when the trade exceeds the risk appetite of the investor, this procedure balances the differing expectations of the counterparties to identify their equilibrium price, accounting for their respective funding costs and risks of default.

### v4

@@ I Introduction
In this essay, foundations for pricing are considered that do not rely on unrealistic assumptions on market completeness or the effectiveness of the trading strategy, properly account for residual model risk, and establish consistency across diverse applications including algorithmic trading, derivative pricing and hedging, and the modelling of margin and capital. In developing these foundations, the entropic risk metric has dual application: as the a priori measure of performance used to optimise the investment strategy; and as the determinant of margin used to protect against default.
@@ III Entropic pricing and hedging
This model is extensive in scope, consistently encompassing arbitrary economic assumptions, predictable and unpredictable funding, complete and incomplete hedge markets, continuous and discontinuous evolution, continuous and discrete hedging, and accounting for the risk appetite of the investor. The case studies in figures 7, LABEL: and 8 demonstrate entropic hedging and pricing for call and digital options over a single investment horizon, and show the bid-offer spread around the mid model that prices the residual risk after hedging. To further develop intuition on the meaning of the hedge and price conditions and their solutions, it is useful to consider boundary cases where they simplify.
@@ III.4 Markovian diffusion
With this assumption, the price model is established on the divisible process for the state vector $x$ and the underlying price vector $q[t,x]$ given as a function of time and state. Using the expression for Itô’s lemma in the price measure, the calibration, hedge and price conditions are transformed into partial differential-integral form. The calibration condition becomes:
@@ III.5 Heston model
The funding price has predictable increment with mean $br\,dt$ where $r$ is the funding rate. The underlying price has increment with mean $q\mu\,dt$ and variance $q^{2}\nu\,dt$ where $\mu$ is the mean rate. The variance has increment with mean $(\bar{\nu}-\nu)\kappa\,dt$ and variance $\nu\sigma^{2}\,dt$ where $\kappa$ is the mean reversion rate, $\bar{\nu}$ is the mean reversion level and $\sigma$ is the volatility of variance. The final parameter $\rho$ is the correlation between the increments of the underlying price and the variance.
Since the funding security has predictable return, the implied funding rate is $r$ and the implied funding spread is $s=0$. In order to satisfy the self-funding condition, the portfolio with underlying weight $\phi$ must have funding weight $-\phi q/b$, and the cumulant generating function in the price measure $\mathbb{E}^{\phi}$ satisfies:
@@
Figure 10: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 1,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 1,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
@@ III.6 Data-driven model
For this data-driven approach, the expectation measure is represented as a uniform distribution on a finite set of states, which could for example be sampled from a generative model trained on historical market data. The calibration, hedge and price conditions are discretised as:
00 0 $=\sum\nolimits_{n}\gamma_{n}(dq_{n}-qr\,dt)$ (74)
00 0 $=\sum\nolimits_{n}\gamma_{n}(dq_{n}-q(r+\alpha s)\,dt)\exp[-\alpha(P_{n}-\delta\cdot dq_{n})]$
$p$ $=-\frac{1}{\alpha}\log\sum\nolimits_{n}\gamma_{n}\exp[-\alpha(P_{n}-\delta\cdot dq_{n})]$
where the price measure is supported on the same set of states but with non-uniform probabilities:
These conditions are accurately and efficiently solved using the Newton-Raphson scheme.
As demonstrated in the examples of figures 9, 10, 11, 12, 13 and 14, the hedge and price conditions can be solved exactly on the assumption of the discrete expectation measure, but the solution is impacted by the choice. In the example with 100 sample points, the support of the measure does not extend beyond the interval (-2.5,2.5), and there can be no option value outside this range of strikes. Increasing the number of points brings the solutions closer to the limit of the standard normal variable used to generate the samples. The support of the expectation measure plays a crucial role in determining hedge and price and also the extent to which unhedged risks impact the spread of prices. Unrealistic assumptions on the support mask these model risks and can hinder their quantification.
There are situations where it is useful to reframe the calibration, hedge and price conditions as optimisation problems. The calibration condition is solved for the unit optimal portfolio $\phi$ via the optimisation problem:
@@
The hedge and price conditions are then solved simultaneously for the hedge portfolio $\delta$ and the initial derivative price $p$ via the optimisation problem:
@@
These expressions are amenable to machine learning methods and are the starting point for a variety of applications in deep hedging. Expressed as optimisation problems, additional features such as frictional trading costs can be included in the program, and large datasets can be accommodated by using universal approximations for the hedging strategy.
The examples of figures 9, 10, 11, 12, 13 and 14 demonstrate the reach of the entropic framework. With no theory to support development beyond the information that the data expresses, data-driven models need a more general foundation that does not depend on unrealistic assumptions such as complete markets, continuous hedging or risk-free funding. Advocates of deep hedging address this by replacing dynamic risk neutralisation with dynamic risk optimisation, using convex risk metrics to assess strategic performance. While they suffer from lack of explainability, data-driven models can uncover complex relationships between economic variables that may be hard to capture theoretically.
@@ V Entropic XVA
Figure 15: In this example, the trader receives the initial derivative price $p$ and deploys additional capital $c$ to guarantee settlement, with margin $m:=c+p\geq 0$. These proceeds are deposited in a margin account with unit price $b>0$. The final derivative price $p+dp$ is settled up to the value $(c+p)(1+db/b)$ of the margin account, with any remainder returned to the trader. Default occurs when the margin account is insufficient to meet the contractual obligation. In the expression for the final settlement, the cap is denoted by the superscript.
@@
Figure 16: Two counterparties deploy capital $c_{1}$ and $c_{2}$ respectively as their guarantee of derivative settlement, with margin $m_{1}:=c_{1}-p\geq 0$ and $m_{2}:=c_{2}+p\geq 0$, and they default on their contractual obligation if this is insufficient. Net of all considerations, the counterparties share the combined proceeds from the margin accounts, and the final return on capital is floored and capped by the respective default events. In the expressions for the final settlement and capital, the floor is denoted by the subscript and the cap is denoted by the superscript.
@@
Figure 17: Compressing internal settlements, the trade is simplified as a funding pool with contributions from both counterparties shared according to the contractual settlement of the derivative security adjusted for default. Using the entropic margin model, these settlements can be encoded in a smart contract for decentralised distribution.
@@
V.1 Margined derivatives
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a derivative trade with initial price $p$. Excess capital is deposited in the margin account with unit price $b>0$. The derivative is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
As shown in figures 17, 17 and 17, net of all considerations the capital returned from the derivative is:
$c_{1}+dc_{1}$ $=c_{1}(1+db/b)+\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$ (85)
$c_{2}+dc_{2}$ $=c_{2}(1+db/b)-\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}$
where the subscript denotes the floor and the superscript denotes the cap. With all the internal settlements compressed, the structure reduces to a funding pool with initial capital $(c_{1}+c_{2})$ deposited in the margin account and final capital $(c_{1}+c_{2})(1+db/b)$ redistributed to the counterparties. Without the derivative settlement, the final capital is distributed according to the return on the initial contribution, with $c_{1}(1+db/b)$ returned to the first counterparty and $c_{2}(1+db/b)$ returned to the second counterparty. Introducing the derivative settlement, this distribution of capital is zero-sum adjusted by the funded derivative return up to the amount in the margin account. Default then occurs when the margin account has insufficient funds to meet the contractual settlement.
With the return on capital thus defined, entropic pricing can be applied. Each counterparty assesses the performance of the derivative security relative to their respective underlying markets with prices $q_{1}$ and $q_{2}$, using their expectation measures $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ and risk aversions $\alpha_{1}$ and $\alpha_{2}$. The hedge conditions:
00 0 $=\mathbb{E}_{1}^{\phi_{1}}[(dq_{1}-q_{1}r_{1}\,dt)\exp[-\alpha_{1}(dc_{1}-\delta_{1}\cdot dq_{1})]]$ (86)
00 0 $=\mathbb{E}_{2}^{\phi_{2}}[(dq_{2}-q_{2}r_{2}\,dt)\exp[-\alpha_{2}(dc_{2}-\delta_{2}\cdot dq_{2})]]$
are solved with the self-funding conditions $\delta_{1}\cdot q_{1}=c_{1}$ and $\delta_{2}\cdot q_{2}=c_{2}$ to generate the hedge portfolios $\delta_{1}$ and $\delta_{2}$, where $\phi_{1}$ and $\phi_{2}$ are the unit optimal portfolios in the underlying markets. The price conditions derive the expressions:
[…]
$\qquad\qquad+\left.\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}\right]$
[…]
$\qquad\qquad-\left.\left(dp-\frac{p}{b}db\right)_{-c_{1}(1+db/b)}^{c_{2}(1+db/b)}\right]$
which are solved for the threshold prices $p_{1}$ and $p_{2}$ of the two counterparties. The trade is viable when the initial price $p$ is in the range $p_{2}\leq p\leq p_{1}$, and is therefore not viable when $p_{1}<p_{2}$. Price is thus constrained to the sector bounded by these viability thresholds, equilibrating to a price within the range.
V.2 Entropic XVA
Embedded in the expressions for the threshold prices are the contributions from XVA.
FVA
Capital for the trade is raised by borrowing in the underlying market and deposited in the margin account. The mismatch between the return on the underlying portfolio and the return on the margin account generates the funding adjustment in the valuation.
CVA
The cap on the potential profit from the funded derivative return generates the credit adjustment in the valuation as a cost to the counterparty.
DVA
The floor on the potential loss from the funded derivative return generates the debit adjustment in the valuation as a benefit to the counterparty.
To understand the XVA contributions, consider the simplified model with predictable funding:
$dq_{1}/q_{1}$ $=r_{)}\,dt$ (88)
$dq_{2}/q_{2}$ $=r_{2}\,dt$
$db/b$ $=r\,dt$
where $r_{1}$ and $r_{2}$ are the funding rates of the underlying securities and $r$ is the funding rate of the margin account. With only one security in each underlying market, the self-funding conditions derive $\delta_{1}=c_{1}/q_{1}$ and $\delta_{2}=c_{2}/q_{2}$ and the threshold prices become:
$p_{1}={}$ $\frac{1}{1+r_{1}\,dt}\mathbb{E}_{1}^{\phi_{1}}[\alpha_{1}]\!\left[P_{-m_{1}(1+r\,dt)}^{m_{2}(1+r\,dt)}\right]$ (89)
$-\frac{m_{1}(r_{1}-r)\,dt}{1+r_{1}\,dt}$
$p_{2}={}$ $\frac{1}{1+r_{2}\,dt}\mathbb{E}_{2}^{\phi_{2}}[-\alpha_{2}]\!\left[P_{-m_{1}(1+r\,dt)}^{m_{2}(1+r\,dt)}\right]$
$+\frac{m_{2}(r_{2}-r)\,dt}{1+r_{2}\,dt}$
where $P:=p+dp$ is the final derivative price and the margins are $m_{1}:=c_{1}-p$ and $m_{2}:=c_{2}+p$ for the two counterparties. The first term is the discounted entropy-adjusted mean in the price measure of the final derivative price, capped and floored by the respective margins to generate the credit and debit valuation adjustments. The second term then adjusts the initial price by the differential between the returns on margin from the underlying security versus the margin account, creating the funding valuation adjustment.
V.3 Entropic margin
The combined proceeds from the margin accounts are shared between the counterparties, adjusting for the funded return on the derivative security, floored and capped so that the returned capital is positive.
The first counterparty defaults when the floor is breached and the second counterparty defaults when the cap is breached. Each counterparty assesses the threshold for default using their expectation measure, and default is avoided when the margin exceeds the maximum liability:
$m_{1}$ $:=c_{1}-p\geq-\operatorname*{ess\,inf}\nolimits_{1}\!\left[\frac{P}{1+db/b}\right]$ (90)
$m_{2}$ $:=c_{2}+p\geq\operatorname*{ess\,sup}\nolimits_{2}\!\left[\frac{P}{1+db/b}\right]$
where $P:=p+dp$ is the contractual settlement of the derivative security. Setting these inequalities to equalities generates the minimum capital required to fully guarantee settlement, under the assumptions of the respective expectation measures.
More generally, the counterparties can exchange certainty of settlement for a reduction in capital. In the entropic margin model, this is parametrised by the default aversion $\beta_{1}$ and $\beta_{2}$ with:
$m_{1}$ $=\left(-\mathbb{E}_{1}[\beta_{1}]\!\left[\frac{P}{1+db/b}\right]\right)^{\!+}$ (91)
$m_{2}$ $=\left(\mathbb{E}_{2}[-\beta_{2}]\!\left[\frac{P}{1+db/b}\right]\right)^{\!+}$
where the case $\beta_{1}=\infty$ means the first counterparty provides full default protection and the case $\beta_{2}=\infty$ means the second counterparty provides full default protection. Entropy-adjusted mean is utilised in this parametrisation as it spans the range of possible values for the discounted derivative settlement.
Trading on margin is the great innovation of finance, but the protections offered to counterparties are often opaque and operationally cumbersome, and occasionally ineffective. With the derivative settlements net of funding and default compressed into a funding pool with guaranteed settlements, and an entropic margin model that codifies the implied level of default protection, entropic pricing can be automated in a smart contract that transparently declares the trade-off between return and default risk. By packaging funding with contractual settlements, this internalises the management of margin and capital and avoids their overheads and inefficiencies, supporting the decentralised distribution of derivative securities.